# Hierarchical hyperbolicity of admissible curve graphs and the boundary of marked strata

### v1

@@ 1. Introduction
There is a storied history to using graphs built from curves on a surface $S$ to understand the mapping class group $\operatorname{Mod}(S)$ of the surface and related objects. The most famous example is Harvey’s curve graph, $\mathscr{C}(S)$, which has a vertex for each isotopy class of essential simple closed curve on an orientable surface and an edge when two curves can be realized disjointly [Har81]. The curve graph is a central object in low-dimensional topology, illuminating the mapping class group [Iva97, MM00], the geometry of Teichmüller space [MM98, Raf05] and the structure of hyperbolic 3-manifolds [Min10, BCM12]. Other examples of this paradigm include the pants graph for understanding the Weil–Petersson metric [Bro03, BF06], the Torelli complex and separating curve graph for studying the Torelli subgroup and the Johnson kernel [FI05, BM04], and the disk graph for examining the handlebody group and Heegaard splittings [Hen20, MS13].
Recently, the first author and Salter have shown that the framed mapping class group plays a important role in understanding moduli spaces of abelian differentials [CS22]. Let $S=S_{g,n}$ be a surface of genus $g$ with $n$ punctures. A framing $\phi$ of $S$ is a trivialization of its tangent bundle, or equivalently (up to isotopy), a non-vanishing vector field. The framed mapping class group $\operatorname{FMod}(S,\phi)$ is the subgroup of $\operatorname{Mod}(S)$ that stabilizes the isotopy class of $\phi$. A natural graph of curves on which $\operatorname{FMod}(S,\phi)$ acts is the admissible curve graph, $\operatorname{\mathscr{C}_{adm}}(S,\phi)$, the subgraph of $\mathscr{C}(S)$ spanned by non-separating curves that have winding number 00 with respect to $\phi$. How closely the relationship between $\operatorname{FMod}(S,\phi)$ and $\operatorname{\mathscr{C}_{adm}}(S,\phi)$ mimics the relationship between $\operatorname{Mod}(S)$ and $\mathscr{C}(S)$ is an open question.
A marquee result on the curve graph $\mathscr{C}(S)$ is Masur and Minsky’s proof of Gromov hyperbolicity [MM98]. This has had far-reaching implications for the coarse geometry of the mapping class group and is a central component of the resolution of the ending lamination conjecture [Min10]. A starting place to understand the relationship between $\operatorname{FMod}(S,\phi)$ and $\operatorname{\mathscr{C}_{adm}}(S,\phi)$ is thus to ask about the geometry of $\operatorname{\mathscr{C}_{adm}}(S,\phi)$. It turns out that the admissible curve graph is not Gromov hyperbolic, but does possess a generalized notion of hyperbolicity.
@@
Hierarchical hyperbolicity was introduced by Behrstock, Hagen, and Sisto to unify similarities between the coarse geometry of mapping class groups, Teichmüller spaces, and right-angled Artin groups [BHS17b]. This framework allows one to understand the geometry of a space by projecting it onto a collection of Gromov hyperbolic spaces. In the case of $\operatorname{\mathscr{C}_{adm}}(S,\phi)$, we use Masur and Minsky’s subsurface projection maps to project $\operatorname{\mathscr{C}_{adm}}(S,\phi)$ on the curve graphs of witnesses — subsurfaces of $S$ that intersect every admissible curve. The lack of Gromov hyperbolicity of $\operatorname{\mathscr{C}_{adm}}(S,\phi)$ emerges from the fact that there exist pairs of disjoint witnesses for $\operatorname{\mathscr{C}_{adm}}(S,\phi)$. Using the hierarchically hyperbolic machinery, these disjoint witnesses induce undistorted product regions in $\operatorname{\mathscr{C}_{adm}}(S,\phi)$ which obstruct Gromov hyperbolicity.
This approach was inspired by work of Vokes, who showed that a wide variety of graphs of curves are similarly hierarchically hyperbolic using their subsurface projection maps to witnesses [Vok22]. Vokes first uses the set of witnesses to build a “model graph” $\mathcal{K}$ that she proves is hierarchically hyperbolic. She then shows that when the graph of curves admits a cobounded action of $\operatorname{Mod}(S)$ it is quasi-isometric to the model graph $\mathcal{K}$. While we can construct Vokes’s hierarchically hyperbolic model graph $\mathcal{K}$ for $\operatorname{\mathscr{C}_{adm}}(S,\phi)$, we cannot employ her quasi-isometry as $\operatorname{\mathscr{C}_{adm}}(S,\phi)$ does not admit an action by all of $\operatorname{Mod}(S)$ and the action of $\operatorname{FMod}(S,\phi)$ on $\mathcal{K}$ is not sufficiently cofinite to adapt her argument. Instead, we must construct a quasi-isometry $\mathcal{K}\to\operatorname{\mathscr{C}_{adm}}(S,\phi)$ by hand.
1.1. Curve graphs for strata
In addition to its topological definition, $\mathscr{C}(S)$ can also be interpreted as encoding the intersection pattern of the boundary of the augmented Teichmüller space $\overline{\mathcal{T}_{g,n}}$ (see §5.2). We can use our analysis of $\operatorname{\mathscr{C}_{adm}}(S,\phi)$ to also understand the geometry of how marked strata of holomorphic abelian differentials meet $\partial\overline{\mathcal{T}_{g,n}}$.
We first recall that an abelian differential is a holomorphic 1-form on a Riemann surface. The moduli space of all abelian differentials of genus $g$ forms a rank $g$ (orbifold) vector bundle $\Omega\mathcal{M}_{g}$ over the usual moduli space of genus $g$ closed Riemann surfaces $\mathcal{M}_{g}$. This bundle is broken into pieces called strata, which parametrize those differentials with a fixed number and order of zeros. Since strata parametrize Riemann surfaces with marked points, and differentials are determined up to scaling by the order and position of their zeros, we may as well think of strata as subvarieties of $\mathcal{M}_{g,n}$, the moduli space of genus $g$ Riemann surfaces with $n$ marked points. Strata are not always connected, but Kontsevich and Zorich classified their connected components [KZ03]: there are always at most $3$, at most one of which is hyperelliptic, in that it consists entirely of hyperelliptic Riemann surfaces with point sets that are invariant under the hyperelliptic involution (preserving orders).
@@ 1.1. Curve graphs for strata
Given a non-hyperelliptic stratum component $\mathcal{H}\subset\mathcal{M}_{g,n}$, let $\mathcal{H}_{\phi}$ be any component of the preimage of $\mathcal{H}$ in $\mathcal{T}_{g,n}$ and consider its closure $\overline{\mathcal{H}_{\phi}}$ in the augmented Teichmüller space. Define a graph $\mathscr{C}(\overline{\mathcal{H}_{\phi}})$ whose vertices are those multicurves $\gamma$ such that $\overline{\mathcal{H}_{\phi}}\cap\mathcal{T}_{g,n}(\gamma)\neq\emptyset$, where $\mathcal{T}_{g,n}(\gamma)$ is the boundary stratum of $\overline{\mathcal{T}_{g,n}}$ in which $\gamma$ is pinched, and whose edges are given by inclusion. This graph is an analogue of the (multi)curve graph for marked strata (Section 5.2, see also Sections 5.3 and 5.4 for a discussion of other natural analogues, all of which end up being quasi-isometric to $\mathscr{C}(\overline{\mathcal{H}_{\phi}})$).
@@
The geometry of this graph is intricately linked to the geometry of the admissible curve graph, and using Theorem A and recent structural results about compactifications of strata [BCG+18, BCG+19] we prove the following.
@@
For any non-hyperelliptic stratum component $\mathcal{H}$ of abelian differentials on a surface of genus $g\geq 5$, the graph $\mathscr{C}(\overline{\mathcal{H}_{\phi}})$ is hierarchically hyperbolic (but not Gromov hyperbolic).
@@
Our restriction to non-hyperelliptic components is because the hyperelliptic ones do not exhibit new phenomena. Indeed, hyperelliptic stratum components are essentially strata of quadratic differentials on the sphere, which are in turn parametrized by their poles and zeros. Thus we can understand compactifications of hyperelliptic stratum components entirely in terms of the Deligne–Mumford compactification of $\mathcal{M}_{0,n}$, and the intersection pattern of the boundary of $\mathcal{T}_{0,n}$ is just the usual curve graph of an $n$-times punctured sphere.
@@
As shown in [CS22, Corollary 1.2], admissible curves are exactly the core curves of cylinders, which give directions to degenerate into the boundary of $\overline{\mathcal{H}_{\phi}}$. One can also construct a partial bordification of ${\mathcal{H}}_{\phi}$ in which only cylinders are allowed to degenerate; the combinatorics of how this space meets $\partial\overline{\mathcal{T}_{g,n}}$ correspond exactly to $\operatorname{\mathscr{C}_{adm}}(S,\phi)$. Thus Theorem A can also be viewed as a statement about the coarse geometry of ${\mathcal{H}}_{\phi}$.
@@
The reason we must restrict to genus $\geq 3$ in Theorem A is because for $g=1,2$ the admissible curve graph is not necessarily connected. The genus restriction to $g\geq 5$ in Theorem B comes from the fact that the main theorem of [CS22] relating $\pi_{1}(\mathcal{H})$ and the framed mapping class group only applies for $g\geq 5$. In Section 5 we give a (slightly circuitous) definition of $\mathscr{C}(\overline{\mathcal{H}_{\phi}})$ that agrees with the one given above for $g\geq 5$ and for which Theorem B holds in genus 3 and 4. In particular, all of the proofs in this paper hold for $g\geq 3$.
Outline of paper
We begin in Section 2 by recalling some basic information on the framed mapping class group and proving fundamental “change-of-coordinate” style lemmas. After these preliminaries, we prove Theorem A in Sections 3 and 4. The former section records Vokes’s construction of a model graph given a collection of witnesses, while in the latter we build a quasi-isometry between the model and $\operatorname{\mathscr{C}_{adm}}(S,\phi)$. In Section 5, we discuss the relevant background on strata and their boundaries and build a trio of analogues for the curve graph. In the final Section 6, we show that the three graphs from Section 5 are all quasi-isometric, and that they are quasi-isometric to a cone-off of $\operatorname{\mathscr{C}_{adm}}(S,\phi)$.
@@
A common theme running throughout this paper is that if one understands the $\operatorname{FMod}(S,\phi)$ action on configurations of curves and subsurfaces well enough, then many surface-topological arguments can be adapted to the framed setting with a little extra care. As such, we prove a number of transitivity results (Propositions 2.9, 6.2, and 6.3) for the $\operatorname{FMod}(S,\phi)$ action that may be of broader interest.

### v2

@@ 1. Introduction
The moduli space $\Omega\mathcal{M}_{g}$ of genus $g$ Abelian differentials forms a bundle over the usual moduli space $\mathcal{M}_{g}$ of genus $g$ Riemann surfaces. This bundle decomposes into strata, subvarieties which parametrize differentials with a given number and order of zeros and which are the ambient theatre for Teichmüller dynamics. The overall structure of strata is still poorly understood, and recent work has been largely guided by the following:
Question 1.1.
How similar are strata and $\mathcal{M}_{g}$?
There has been a great deal of success constructing compactifications of strata akin to the Deligne–Mumford compactification of $\mathcal{M}_{g}$ [EMZ03, BCG+18, BCG+19]. The structure of these boundaries can then be used to compute constants of dynamical interest [EMZ03], perform intersection theory on strata [CMSZ20], and compute their Euler characteristics [CMZ22], among many other things.
Another version of Question 1.1 deals with their fundamental groups. Recall that $\mathcal{M}_{g}$ is an (orbifold) $K(\pi,1)$ for the usual mapping class group $\operatorname{Mod}(S)$, the group of homeomorphisms of the surface up to homotopy. By analogy, Kontsevich predicted that each connected component of a stratum should be a $K(\pi,1)$ for “some mapping class group” [KZ]. In [CS22], the first author and Salter showed that the fundamental groups of strata are closely related to framed mapping class groups $\operatorname{FMod}(S,\phi)$, the stabilizers inside $\operatorname{Mod}(S)$ of trivializations $\phi:TS\cong S\times\mathbb{R}^{2}$ (see §2 for a formal definition). Apisa, Bainbridge, and Wang subsequently showed that certain strata of twisted 1-forms are $K(\pi,1)$’s for framed mapping class groups [ABW23]. A group-theoretic analogue of Question 1.1 is thus:
Question 1.2.
How similar are $\operatorname{FMod}(S,\phi)$ and $\operatorname{Mod}(S)$?
1.1. Curve graphs and strata
This paper initiates the study of Questions 1.1 and 1.2 from the coarse-geometric perspective by analyzing the geometry of certain curve graphs.
The classical curve graph $\mathscr{C}(S)$ has a vertex for each isotopy class of essential simple closed curve on an orientable surface and an edge when two curves can be realized disjointly [Har81]. In addition to this topological interpretation, this graph also plays the role of (the 1-skeleton of) a Tits building for Teichmüller space $\mathcal{T}_{g}$, recording the incidences of top-dimensional boundary strata of the augmented Teichmüller space, a certain bordification of $\mathcal{T}_{g}$ that “lifts” the Deligne– Mumford compactification of $\mathcal{M}_{g}$ (see §5.2).
Masur and Minsky famously proved that $\mathscr{C}(S)$ is Gromov hyperbolic [MM98]. This marquee result has far-reaching implications for the coarse geometry of the mapping class group [Iva97, MM00], the geometry of Teichmüller space [MM98, Raf05], and the structure of hyperbolic 3-manifolds [Min10, BCM12]. More generally, the geometry of curve graphs has proven useful in a variety of settings; examples of this paradigm include relationships between the pants graph/the Weil–Petersson metric on $\mathcal{T}_{g}$ [Bro03, BF06], the Torelli complex and separating curve graph/the Torelli subgroup and the Johnson kernel [FI05, BM04], and the disk graph/the handlebody group and Heegaard splittings [Hen20, MS13].
As a first step towards Question 1.2, we study a topological analogue of $\mathscr{C}(S)$ that takes the framing into account. Any framing $\phi:TS\cong S\times\mathbb{R}^{2}$ can be used to measure the winding number of a smooth, oriented curve in $S$ by lifting the curve to $TS$ via its tangent vector, projecting to the second coordinate, then measuring the winding number of the image about $0\in\mathbb{R}^{2}$. A simple closed curve on $S$ is admissible for $\phi$ if it is nonseparating and has zero winding number, and the admissible curve graph $\operatorname{\mathscr{C}_{adm}}(S,\phi)$ is the subgraph of $\mathscr{C}(S)$ spanned by admissible curves.
The framed mapping class group $\operatorname{FMod}(S,\phi)$ preserves the winding number of every curve, hence acts on $\mathscr{C}(S)$ with infinitely many orbits of vertices. In contrast, $\operatorname{FMod}(S,\phi)$ acts on $\operatorname{\mathscr{C}_{adm}}(S,\phi)$ with finitely many orbits of vertices and edges (Proposition 2.9), indicating that the admissible curve graph is better adapted to study $\operatorname{FMod}(S,\phi)$.
Our first main result is that the admissible curve graph is not Gromov hyperbolic, but does possess a generalized notion of hyperbolicity.
@@ 1.1. Curve graphs and strata
Hierarchical hyperbolicity was introduced by Behrstock, Hagen, and Sisto to unify similarities between the coarse geometry of mapping class groups, Teichmüller spaces, and right-angled Artin groups [BHS17b]. Briefly, this framework allows one to understand the geometry of a space by projecting it onto a collection of Gromov hyperbolic spaces. The presence of “orthogonal” projections leads to quasi-isometrically embedded flats, hence a failure of Gromov hyperbolicity.
We can also define a geometric analogue of $\mathscr{C}(S)$ that captures the intersection pattern of the boundary of a marked stratum. More precisely, since holomorphic differentials are determined up to scaling by the order and position of their zeros, any stratum component $\mathcal{H}\subset\Omega^{1}\mathcal{M}_{g}$ is an (orbifold) $\mathbb{C}^{*}$-bundle over a subvariety of $\mathcal{M}_{g,n}$, the moduli space of genus $g$ Riemann surfaces with $n$ marked points. Let us conflate $\mathcal{H}$ with this subvariety.
@@
Take any non-hyperelliptic stratum component $\mathcal{H}\subset\mathcal{M}_{g,n}$, let $\mathcal{H}_{\phi}$ be any component of the preimage of $\mathcal{H}$ in $\mathcal{T}_{g,n}$, and consider its closure $\overline{\mathcal{H}_{\phi}}$ in the augmented Teichmüller space $\overline{\mathcal{T}_{g,n}}$. Define a graph $\mathscr{C}(\overline{\mathcal{H}_{\phi}})$ whose vertices are those multicurves $\gamma$ such that $\overline{\mathcal{H}_{\phi}}\cap\mathcal{T}_{g,n}(\gamma)\neq\emptyset$, where $\mathcal{T}_{g,n}(\gamma)$ is the boundary stratum of $\overline{\mathcal{T}_{g,n}}$ in which $\gamma$ is pinched, and whose edges are given by inclusion. The intricate structure of the boundary of $\overline{\mathcal{H}_{\phi}}$ means there are other natural ways to define this graph (see Sections 5.3 and 5.4), but they all turn out to be quasi-isometric to $\mathscr{C}(\overline{\mathcal{H}_{\phi}})$.
@@
The geometry of $\mathscr{C}(\overline{\mathcal{H}_{\phi}})$ is closely linked to that of $\operatorname{\mathscr{C}_{adm}}(S,\phi)$, and using Theorem A plus structural results about compactifications of strata [BCG+18, BCG+19], we prove:
@@
For any non-hyperelliptic stratum component $\mathcal{H}\subset\Omega^{1}\mathcal{M}_{g}$ with $g\geq 5$, the graph $\mathscr{C}(\overline{\mathcal{H}_{\phi}})$ is hierarchically hyperbolic (but not Gromov hyperbolic).
@@
As shown in [CS22, Corollary 1.2], admissible curves are exactly the core curves of cylinders on surfaces in ${\mathcal{H}}_{\phi}$. One can also construct a partial bordification of ${\mathcal{H}}_{\phi}$ in which only cylinders are allowed to degenerate; the combinatorics of how this space meets $\partial\overline{\mathcal{T}_{g,n}}$ then correspond to $\operatorname{\mathscr{C}_{adm}}(S,\phi)$. Thus Theorem A can also be interpreted as a statement about the coarse geometry of ${\mathcal{H}}_{\phi}$.
@@
Our restriction to non-hyperelliptic components is because the hyperelliptic ones do not exhibit new phenomena. Indeed, hyperelliptic stratum components are essentially strata of quadratic differentials on $\mathbb{CP}^{1}$, which are in turn parametrized by their poles and zeros. Thus we can understand compactifications of hyperelliptic stratum components entirely in terms of the Deligne–Mumford compactification of $\mathcal{M}_{0,n}$.
@@
The restriction to $g\geq 3$ in Theorem A is because for $g=1,2$ the admissible curve graph is not necessarily connected. The restriction to $g\geq 5$ in Theorem B comes from the fact that the main theorem of [CS22] relating $\pi_{1}(\mathcal{H})$ and $\operatorname{FMod}(S,\phi)$ only applies for $g\geq 5$. In Section 5 we give a (slightly circuitous) definition of $\mathscr{C}(\overline{\mathcal{H}_{\phi}})$ that agrees with the one given above for $g\geq 5$ and for which Theorem B holds in genus 3 and 4. In particular, all of the proofs in this paper hold for $g\geq 3$.
Curve graph techniques have been used successfully to study certain $\mathsf{GL}_{2}\mathbb{R}$–invariant subvarieties of $\Omega^{1}\mathcal{M}_{g}$: [Tan21] proved that Veech groups are undistorted in $\operatorname{Mod}(S)$, [RS09] proved a similar result for covering constructions, and [AHW24] used curve graphs to study the geometry of totally geodesic subvarieties of Teichmüller space. It is our hope that the tools developed in this paper will yield insights into both the intrinsic and extrinsic geometry of framed mapping class groups and strata. For example, we ask:
Question 1.6.
Is $\operatorname{FMod}(S,\phi)$ distorted in $\operatorname{Mod}(S)$? Are strata distorted in $\mathcal{M}_{g,n}$?
1.2. Outline of proof and paper
To prove Theorems A and B, we need to exhibit projections from $\operatorname{\mathscr{C}_{adm}}(S,\phi)$ and $\mathscr{C}(\overline{\mathcal{H}_{\phi}})$ to Gromov hyperbolic spaces. In both settings, we use Masur and Minsky’s subsurface projection maps to the curve graphs of witnesses — subsurfaces of $S$ that intersect every admissible curve. This approach was inspired by work of Vokes, who showed that a wide variety of graphs of curves are hierarchically hyperbolic using their subsurface projection maps to witnesses [Vok22]. Vokes first uses the set of witnesses to build a hierarchically hyperbolic “model graph” $\mathcal{K}$, then shows that if the graph of curves admits a cobounded action of $\operatorname{Mod}(S)$ then it is quasi-isometric to $\mathcal{K}$.
To prove Theorem A, we construct a hierarchically hyperbolic model $\mathcal{K}$ for $\operatorname{\mathscr{C}_{adm}}(S,\phi)$ la Vokes (Section 3). However, we cannot employ her quasi-isometry as $\operatorname{\mathscr{C}_{adm}}(S,\phi)$ does not admit an action by all of $\operatorname{Mod}(S)$ and the action of $\operatorname{FMod}(S,\phi)$ on $\mathcal{K}$ is not sufficiently cofinite to adapt her argument. Instead, we construct a novel quasi-isometry $\mathcal{K}\to\operatorname{\mathscr{C}_{adm}}(S,\phi)$ via the graph $\mathcal{G}$ of genus separating curves (Section 4). The graph $\mathcal{G}$ can be quasi-isometrically realized as a “blow-up” of $\mathcal{K}$, while $\operatorname{\mathscr{C}_{adm}}(S,\phi)$ is quasi-isometric to a “cone-off” of $\mathcal{G}$. To build the map $\mathcal{K}\to\operatorname{\mathscr{C}_{adm}}(S,\phi)$, we show that the blown-up subsets from $\mathcal{K}\to\mathcal{G}$ coarsely match the coned-off subsets from $\mathcal{G}\to\operatorname{\mathscr{C}_{adm}}(S,\phi)$. This step requires some fairly delicate computations with curves on surfaces.
Theorem B follows by constructing a quasi-isometric model for $\mathscr{C}(\overline{\mathcal{H}_{\phi}})$ entirely in terms of framing data. This requires unpacking some of the finer structure of the boundary, as developed in [BCG+19], and giving topological interpretations to many of the objects involved. These steps are accomplished in Section 5. In this section, we also build a trio of graphs whose definitions interpolate between the structure of $\partial\overline{\mathcal{H}_{\phi}}$ and framing data.
In the final Section 6, we show that the three graphs from Section 5 are all quasi-isometric, and that they are quasi-isometric to a Vokes model graph $\overline{\mathcal{K}}$. Again, there is not sufficient transitivity to apply Vokes’s methods, and the construction of a quasi-isometry is quite subtle. The graph $\overline{\mathcal{K}}$ is an $\operatorname{FMod}(S,\phi)$–equivariant cone-off of the model $\mathcal{K}$ for $\operatorname{\mathscr{C}_{adm}}(S,\phi)$, and the inclusion $\operatorname{\mathscr{C}_{adm}}(S,\phi)\hookrightarrow\mathscr{C}(\overline{\mathcal{H}_{\phi}})$ is also an equivariant cone-off. As in the case of Theorem A, the main difficulty is then showing that these two cone-offs coarsely match.
@@ 1.2. Outline of proof and paper
A common theme running throughout this paper is that if one understands the $\operatorname{FMod}(S,\phi)$ action on configurations of curves and subsurfaces well enough, then many surface-topological arguments can be adapted to the framed setting with a little extra care and effort. As such, we prove a number of transitivity results (Propositions 2.9, 6.2, and 6.3) for the $\operatorname{FMod}(S,\phi)$ action that may be of broader interest.